Session: Feb 1999 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (T. Venturi at Mon Feb 15 08:16:08 MET 1999) Success
The experiment went well. No apparent problems to be reported.

Onsala (A. Polatidis at Mon Feb 15 09:48:38 MET 1999) Failure
No observations due to strong winds, higher than the limit for telescope operation.

Torun (K.Borkowski at Mon Feb 15 13:21:12 MET 1999) Success
No known problems.

Noto (C.Trigilio at Tue Feb 16 18:25:10 MET 1999) Success

Westerbork (Tony Foley at Wed Feb 17 13:50:55 MET 1999) Success
No known problems

Jodrell (Alastair Gunn at Fri Feb 19 14:12:06 MET 1999) Minor failures
All scans started 20 seconds late due to lengthy preob. No other known problems.

Effelsberg (dgraham at Mon Feb 22 15:22:29 MET 1999) Minor failures
Stopped for snow 0315 - 0510

Calibration (Jean\-Francois Desmurs at Wed Feb 24 11:54:27 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: v003c.antab. Read the comments in the file to get further information.